# Learning to Intervene on Concept Bottlenecks

### v2

@@ Abstract
While traditional deep learning models often lack interpretability, concept bottleneck models (CBMs) provide inherent explanations via their concept representations. Specifically, they allow users to perform interventional interactions on these concepts by updating the concept values and thus correcting the predictive output of the model. Traditionally, however, these interventions are applied to the model only once and discarded afterward. To rectify this, we present concept bottleneck memory models (CB2M), an extension to CBMs. Specifically, a CB2M learns to generalize interventions to appropriate novel situations via a two-fold memory with which it can learn to detect mistakes and to reapply previous interventions. In this way, a CB2M learns to automatically improve model performance from a few initially obtained interventions. If no prior human interventions are available, a CB2M can detect potential mistakes of the CBM bottleneck and request targeted interventions. In our experimental evaluations on challenging scenarios like handling distribution shifts and confounded training data, we illustrate that CB2M are able to successfully generalize interventions to unseen data and can indeed identify wrongly inferred concepts. Overall, our results show that CB2M is a great tool for users to provide interactive feedback on CBMs, e.g., by guiding a user’s interaction and requiring fewer interventions.
@@ 1 Introduction
Deep learning models often represent black-box models that make it difficult for human users to understand their decision processes Adadi & Berrada (2018); Cambria et al. (2023); Saeed & Omlin (2023) and interact with these Schramowski et al. (2020); Teso et al. (2023). One recent branch within eXplainable Artificial Intelligence (XAI) focuses on the potential of so-called concept bottleneck models (CBMs) Koh et al. (2020); Stammer et al. (2021) to tackle such issues.
@@
These are designed to be inherently human-interpretable. They perform inference (e.g., for bird image classification, cf. Fig. 1 (top)), by transforming the initial raw input into a set of human-understandable concepts (e.g., wing shape and color) with a bottleneck network and provide a final task prediction based on the activation of these concepts with a predictor network. The concept activations thereby serve as an inherent explanation of the model’s decision Teso et al. (2023). Arguably even more valuable, these concept activations can be used as a means for humans to perform targeted interactions, e.g., for querying further explanations Abid et al. (2022) or to correct the model’s concept prediction Koh et al. (2020).
A recent surge of research has focused on the benefits of leveraging interactions in AI models in general Ouyang et al. (2022); Miller (2019), and also CBMs in particular Teso et al. (2023). Multiple such approaches focus on leveraging interactions for mitigating errors of the predictor network Bontempelli et al. (2021); Stammer et al. (2021). However, little work has focused on mitigating potential errors of the initial bottleneck network. Moreover, although interventional interactions on a CBM’s concept activations are a natural tool for this purpose, they have received little attention since their introduction by Koh et al. Koh et al. (2020).
One likely reason for this is that interventions according to Koh et al. (2020) represent a once-use tool for updating model performance by adding human-provided concept labels to an increasing number of randomly selected concepts. For sustainably improving a model’s performance, however, this approach is inefficient and potentially demands a large number of repeated user interactions, where providing such repeated feedback has been identified to lead to a loss in focus of human users Amershi et al. (2014).
In this work, we therefore argue to harvest the rich information present in previously collected interventions in a multi-use approach. Specifically, let us suppose a user corrects a model’s inferred concepts through a targeted intervention. In that case, the intervention carries information on where the model did not perform well. This can be used to improve prediction in similar future situations (Fig. 1 bottom). In this context, we therefore introduce Concept Bottleneck Memory Models (CB2M) as a novel and model-agnostic extension to CBMs. CB2M are based on adding a two-fold memory of interventions to the CBM architecture, which allows to keep track of previous model mistakes as well as previously applied interventions. This memory enables a CB2M to reapply interventions when the CBM repeats mistakes and thus automatically corrects them without the need for additional human feedback. This ultimately allows to overcome the issue of one-time interventions of standard CBMs and enables the model to learn more effectively from provided human feedback. Overall, however, human feedback can be unavailable, and obtaining it is costly. CB2M mitigates this issue by its ability to detect potential model mistakes prior to initial human feedback where its memory module can be used to specifically select data points for human inspection, and thus guide human feedback to where it is really needed.
@@
Figure 1: Reusing a CBM intervention can correct model mistakes for multiple examples. Top: CBMs generate a human interpretable concept representation to solve the final task. Human users can correct these concept predictions via targeted interventions (blue) influencing the final prediction. Bottom: Human interventions hold valuable information that can be reused in suitable situations to automatically correct model errors without further human interactions.
@@
We illustrate the full potential of CB2M in our experimental evaluations on several challenging tasks, such as handling distribution shifts and confounding factors across several datasets. In summary, we make the following contributions: (i) We identify the potential of extracting generalizable knowledge from human interventions as a means of correcting concept bottleneck models. (ii) We introduce CB2M, a flexible extension to any CBM-like architecture for handling such interactive interventions. (iii) Our experimental evaluations show that CB2M can truly learn from interventions by generalizing them to previously unseen examples. (iv) We further show that CB2M are also able to detect model mistakes without the need for initial human knowledge and thus allow to query a user for targeted interventions.
@@
The rest of the paper proceeds as follows: In Sec. 2, we provide a brief background followed by the introduction of CB2M. We present our experiments evaluations in Sec. 3. Afterwards, we relate CB2M to other work in Sec. 4 before concluding the paper together with potential future research directions in Sec. 5.
2 Concept Bottleneck Memory Models (CB2M)
@@
Let us first introduce the background notations on CBMs and interventions before introducing CB2M for detecting model mistakes and generalizing interventions to novel, unseen examples.
@@ 2.1 Background
A CBM which solves the task of transforming inputs $\mathcal{X}$ to outputs $\mathcal{Y}$ consists of two parts. The bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathcal{X}$ into its concept representation $c$. Afterward, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output $y\in\mathcal{Y}$. The correct, i.e., ground-truth values for $c$ and $y$ are written as $c^{*}$ and $y^{*}$, respectively. We refer to overall model (task) accuracy as $\text{Acc}_{f}$ and to concept accuracy as $\text{Acc}_{g}$.
Human interactions with the concept representations are called interventions. An intervention $i\in\mathcal{I}$ is a set of tuples $i=\{(c^{\prime}_{j},j)|j\in\mathcal{J}_{i}\}$, with updated concept values $c^{\prime}_{j}$ and concept indices $j$. $\mathcal{J}_{i}$ is the set of all indices for intervention $i$. Applying an intervention to a sample $x$ overwrites the predicted concept values with those of the intervention, which we denote as $x|i$.
As CBMs consist of two processing modules, the bottleneck and predictor networks, errors can occur in either, with different consequences on how to handle these Bontempelli et al. (2021). If the bottleneck makes an error, this error will most likely also negatively influence the predictor. On the other hand, it is also possible that the predictor makes a wrong final prediction despite having received a correct concept representation. In the latter case, the concept space is either insufficient to solve the task, or the predictor network is susceptible to, e.g., some spurious correlations. Where other works have investigated handling an insufficient concept space through additional (unsupervised) concepts Sawada & Nakamura (2022), or correcting a predictor with spurious correlations Stammer et al. (2021) CB2M focuses on mitigating errors that originate from the bottleneck model. This is achieved by utilizing interventions on the concept space. Let us discuss this in more detail in the following.
@@ 2.2 Concept Bottleneck Memory Models
Figure 2: Overview of CB2M to detect mistakes or generalize interventions. A vanilla CBM (grey) is extended with a two-fold memory (orange and green). The memory compares encodings of new samples to known mistakes to (i) detect model errors or (ii) automatically correct the model via reuse of interventions.
@@
Let us now introduce Concept Bottleneck Memory Models (CB2M) as a flexible extension for any CBM architecture. The bottleneck and predictor networks of the CBM remain unchanged but are extended by a two-fold memory module $\mathcal{M}$ which consists of a mistake memory $\mathcal{M}^{m}$ coupled with a intervention memory $\mathcal{M}^{i}$. The mistake memory operates on encodings $x_{e}$, i.e., the input of the last layer of the bottleneck network. It measures the similarity between two data points $x$ and $x^{\prime}$ via the euclidean distance of their encodings, $d(x_{e},x^{\prime}_{e})=\|x_{e}-x^{\prime}_{e}\|$. The intervention memory directly keeps track of known interventions and associates them to elements of the mistake memory, meaning that the memorized intervention $i$ can be used to correct the memorized mistake of $x_{e}$. We denote an associated encoding and intervention as $\alpha(x_{e},i)$. Overall, this joint memory can be used to detect model mistakes (orange in Fig. 2) or enable automatic reuse of memorized interventions (green in Fig. 2), which we explain in detail in the following paragraphs.
@@
By extending the vanilla CBM with a memory, CB2M can be used for two distinct tasks (cf. Fig. 2): (i) detecting potential model mistakes and (ii) generalizing interventions to new examples. Besides the general advantage of knowing when an AI model has made an incorrect prediction, this knowledge is even more relevant for CBMs as this can be utilized to focus human attention toward obtaining a targeted intervention. Thus, with the ability to handle task (i) CB2M is especially relevant when humans want to provide interventional feedback to a CBM. Furthermore, after humans have intervened on a CBM, they have, in fact, provided valuable knowledge also for future situations. Thus, an applied intervention carries information on how to correct a particular model mistake. We claim that this information should not be discarded as in the original work of Koh et al. Koh et al. (2020), but be reused when similar mistakes occur again. This is where task (ii) of CB2M comes into play.
Detecting Wrongly Classified Instances.
@@
Intuitively, if a data point is similar to other examples where the model has made mistakes, the model will more likely repeat these mistakes on the new data point. Therefore, we utilize the mistake memory $M_{m}$ to keep track of previous mistakes (cf. Alg. 1 for pseudo-code). First, the memory is filled with encodings of datapoints, for which the model did not initially generate the correct output and for which the concept accuracy is smaller than a threshold $t_{a}\in[0,1]$. This leads to: $\mathcal{M}^{m}=\{x_{e}:f(g(x))\neq y^{*}\wedge Acc_{g}(x)<t_{a}\}$. For a new unseen instance $\hat{x}$, we then compare its encoding $\hat{x}_{e}$ with the mistakes in the memory $\mathcal{M}^{m}$ (Alg. 1, lines 4 - 8). If we find $k$ mistakes with a distance to $\hat{x}_{e}$ smaller than $t_{d}$, we consider a model to be making a known mistake. Formally, we thus predict a model mistake for a new unseen instance $\hat{x}$ if:

### v3

@@ Abstract
While deep learning models often lack interpretability, concept bottleneck models (CBMs) provide inherent explanations via their concept representations. Moreover, they allow users to perform interventional interactions on these concepts by updating the concept values and thus correcting the predictive output of the model. Up to this point, these interventions were typically applied to the model just once and then discarded. To rectify this, we present concept bottleneck memory models (CB2Ms), which keep a memory of past interventions. Specifically, CB2Ms leverage a two-fold memory to generalize interventions to appropriate novel situations, enabling the model to identify errors and reapply previous interventions. This way, a CB2M learns to automatically improve model performance from a few initially obtained interventions. If no prior human interventions are available, a CB2M can detect potential mistakes of the CBM bottleneck and request targeted interventions. Our experimental evaluations on challenging scenarios like handling distribution shifts and confounded data demonstrate that CB2Ms are able to successfully generalize interventions to unseen data and can indeed identify wrongly inferred concepts. Hence, CB2Ms are a valuable tool for users to provide interactive feedback on CBMs, by guiding a user’s interaction and requiring fewer interventions.
Machine Learning, ICML, CBM, Interventions, Interactive ML
@@ 1 Introduction
Deep learning models are often deemed black-box models that make it difficult for human users to understand their decision processes (Adadi & Berrada, 2018; Cambria et al., 2023; Saeed & Omlin, 2023) and interact with them (Schramowski et al., 2020; Teso et al., 2023). To address these issues, one recent branch within explainable artificial intelligence focuses on the potential of concept bottleneck models (CBMs) (Koh et al., 2020; Stammer et al., 2021). These are designed to be partially interpretable and perform inference (such as bird image classification cf. Fig. 1 top) by transforming the initial raw input into a set of human-understandable concepts (e.g., wing shape or color) with a bottleneck network. Subsequently, a predictor network provides a final task prediction based on the activation of these concepts. These concept activations serve as an inherent explanation of the model’s decision (Teso et al., 2023). Arguably even more valuable, these activations can be used by humans to perform interventional interactions, e.g., for querying further explanations (Abid et al., 2022) or correcting concept predictions (Koh et al., 2020).
@@
Figure 1: Reusing a CBM intervention can correct model mistakes for multiple examples. Top: CBMs generate a human interpretable concept representation via bottleneck ($g$) to solve the final task with a predictor ($f$). Human users can correct these concept predictions via targeted interventions (blue) influencing the final prediction. Bottom: Human interventions hold valuable information reusable in the right situations to automatically correct model errors without further human interactions.
In fact, a recent surge of research has focused on the benefits of leveraging interactions in AI models in general (Ouyang et al., 2022; Miller, 2019), and also CBMs in particular (Teso et al., 2023). Multiple such approaches focus on leveraging interactions for mitigating errors of the predictor network (Bontempelli et al., 2021; Stammer et al., 2021). So far, little work has focused on mitigating errors in the initial bottleneck network. Moreover, although interventional interactions on a CBM’s concept activations are a natural tool for this purpose, they have received little attention since their introduction by Koh et al. (2020). One likely reason for this is that interventions according to (Koh et al., 2020) represent a singular-use tool for updating model performance by adding human-provided concept labels to an increasing number of randomly selected concepts. For sustainably improving a model’s performance, however, this approach is inefficient and potentially demands a large number of repetitive user interactions. Providing such repeated feedback has been identified to lead to a loss in focus of human users (Amershi et al., 2014) if not infeasible at all.
In this work, we therefore argue to harvest the rich information present in previously collected interventions in a multi-use approach. Specifically, let us suppose a user corrects a model’s inferred concepts through a targeted intervention. In that case, the intervention carries information on where the model did not perform well. As shown in Fig. 1 bottom, this information can be used to improve predictions in similar future situations. In this context, we introduce Concept Bottleneck Memory Models (CB2Ms) as a novel and flexible extension to CBMs. CB2Ms are based on adding a two-fold memory of interventions to the CBM architecture, which allows to keep track of previous model mistakes as well as previously applied interventions. This memory enables two important properties for improved interactive concept learning. Specifically, a CB2M can (1) reapply interventions when the base CBM repeats previous mistakes. It thereby automatically corrects these mistakes without the need for additional human feedback. Overall, human feedback may, however, not always be readily available, and obtaining it can be costly. CB2M thus mitigates this issue by (2) its ability to detect potential model mistakes prior to initial human feedback. Its memory module can be used to select data points for human inspection, and thus guide human feedback to where it is really needed. Thus ultimately, CB2Ms allow to overcome the issue of one-time interventions of standard CBMs and enables the model to learn more effectively from targeted human feedback.
@@
We illustrate the full potential of CB2M in our experimental evaluations on several challenging tasks, such as handling distribution shifts and confounding factors across several datasets. In summary, we make the following contributions: (i) We identify the potential of extracting generalizable knowledge from human interventions as a means of correcting concept bottleneck models. (ii) We introduce CB2M, a flexible extension to CBM-like architectures for handling such interactive interventions. (iii) Our experimental evaluations show that CB2Ms can truly learn from interventions by generalizing them to unseen examples. (iv) We further show that CB2Ms are also able to detect model mistakes without the need for initial human knowledge and thus allow to query a user for targeted interventions 111code is available at: https://github.com/ml-research/CB2M.
@@
We proceed as follows: Sec. 2, provides a brief background followed by the introduction of CB2M. The experiment evaluations are presented in Sec. 3. Afterwards, we relate CB2M to other work in Sec. 4 before concluding the paper together with potential future research directions in Sec. 5.
Figure 2: Overview of CB2M to detect mistakes or generalize interventions. A vanilla CBM (grey), consisting of bottleneck ($g$) and predictor ($f$), is extended with a two-fold memory (orange and green). The memory compares encodings of new samples to known mistakes to (i) detect model errors or (ii) automatically correct the model via reuse of interventions.
2 Concept Bottleneck Memory Models (CB2Ms)
@@
Let us first introduce the background notations on CBMs and interventions before presenting CB2Ms to improve interactive concept learning via detecting of model mistakes and generalizing of interventions to novel, unseen examples.
@@ 2.1 Background
A CBM that solves the task of transforming inputs $\mathcal{X}$ to outputs $\mathcal{Y}$ consists of two parts. The bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathcal{X}$ into its concept representation $c$. Afterward, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output $y\in\mathcal{Y}$. The ground-truth values for $c$ and $y$ are written as $c^{*}$ and $y^{*}$, respectively. We refer to overall model (task) accuracy as $\text{Acc}_{f}$ and to concept accuracy as $\text{Acc}_{g}$. Human interactions with concept representations are called interventions. An intervention $i\in\mathcal{I}$ in the context of CBMs is a set of tuples $i=\{(c^{\prime}_{j},j)|j\in\mathcal{J}_{i}\}$, with updated concept values $c^{\prime}_{j}$ and concept indices $j$. $\mathcal{J}_{i}$ is the set of all indices for intervention $i$. Applying an intervention to a sample $x$ overwrites the predicted concept values with those of the intervention, which we denote as $x|_{i}$. This naturally builds on the assumption that correcting concepts via interventions will improve the output of the CBM which necessitates training an independent CBM in contrast to jointly trained CBMs (Koh et al., 2020; Shin et al., 2023).
As CBMs consist of two processing modules, the bottleneck and predictor networks, errors can occur in either, with different consequences on how to handle these (Bontempelli et al., 2021). If the bottleneck makes an error, this error will most likely also negatively influence the predictor. On the other hand, it is also possible that the predictor makes a wrong final prediction despite having received a correct concept representation. In the latter case, the concept space is either insufficient to solve the task, or the predictor network is susceptible to, e.g., some spurious correlations. Where other works have investigated handling an insufficient concept space through additional (unsupervised) concepts (Sawada & Nakamura, 2022), or correcting a predictor with spurious correlations (Stammer et al., 2021) CB2M, focuses on mitigating errors that originate from the bottleneck model. This is achieved by utilizing interventions in the concept space. Let us now discuss this in more detail.
@@ 2.2 Concept Bottleneck Memory Models
Let us now introduce Concept Bottleneck Memory Models (CB2Ms) as a flexible extension to CBM architectures. The bottleneck and predictor networks of the CBM remain unchanged but are extended by a two-fold memory module $\mathcal{M}$ which consists of a mistake memory $\mathcal{M}^{M}$ coupled with an intervention memory $\mathcal{M}^{I}$. The mistake memory operates on encodings $x_{e}$, i.e., the input of the last layer of the bottleneck network. It measures the similarity between two data points $x$ and $x^{\prime}$, i.e., via the euclidean distance of their encodings, $d(x_{e},x^{\prime}_{e})=\|x_{e}-x^{\prime}_{e}\|$. The intervention memory directly keeps track of known interventions and associates them to elements of the mistake memory, meaning that the memorized intervention $i$ can be used to correct the memorized mistake of $x_{e}$. We denote an associated encoding and intervention as $\alpha(x_{e},i)$.
Overall, the mistake memory can be used to detect model mistakes (orange in Fig. 2) or, together with the intervention memory, enable the automatic reuse of memorized interventions (green in Fig. 2), which we explain in detail in the following paragraphs. Importantly, the character of this memory is independent of the overall CB2M framework. It can be constructed in a differentiable manner, e.g., with neural nearest neighbors (Plötz & Roth, 2018) or, based on traditional nearest neighbor algorithms.
@@
By extending the vanilla CBM with a memory, CB2M can be used for two distinct tasks (cf. Fig. 2): (i) detecting potential model mistakes and (ii) generalizing interventions to new examples. Besides the general advantage of knowing when an AI model has made an incorrect prediction, this knowledge is even more relevant for CBMs as human users can be queried for interventions in a targeted fashion. Thus, the ability to handle task (i) via CB2M is especially relevant when humans want to provide interventional feedback to a CBM. Furthermore, after humans have intervened on a CBM, they have, in fact, provided valuable knowledge for future situations. We claim that this information should not be discarded as in the original work of Koh et al. (2020), but be reused when similar mistakes occur again. This is where task (ii) of CB2M comes into play.
@@
Detecting Wrongly Classified Instances. Intuitively, if a data point is similar to other examples where the model made mistakes, the model will more likely repeat these mistakes on the novel data point. Therefore, in CB2Ms the mistake memory $M_{m}$ is utilized to keep track of previous mistakes (cf. Alg. 1 in the appendix). First, the memory is filled with encodings of datapoints, for which the model did not initially generate the correct output and for which the concept accuracy is smaller than a threshold $t_{a}\in[0,1]$. This leads to: $\mathcal{M}^{M}=\{x_{e}:f(g(x))\neq y^{*}\wedge\mathrm{~{}Acc}_{g}(x)<t_{a}\}$. For a new unseen instance $\hat{x}$, we then compare its encoding $\hat{x}_{e}$ with the mistakes in the memory $\mathcal{M}^{M}$. If we find $k$ mistakes with a distance to $\hat{x}_{e}$ smaller than $t_{d}$, we consider a model to be making a known mistake. Formally, we predict a model mistake for a new unseen instance $\hat{x}$ if: